The area of a square is 45 more square inches than its perimeter. Find the length of each side of the square.
Answer by usyim88hk(158)

let x be one side of the square:
x^2 is the area of the square
4x is the perimeter of the square
x^2 = 4x + 45 (Area is equal to the perimeter plus 45)
x^2-4x-45 = 4x+45-4x-45
x^2-4x-45 = 0 (now factor)
(x-9)(x+5) = 0
x = 9 or -5
Since length cannot be a negative number, -5 is not the answer, so the answer is 9
-------------------------------------------------------------------------------
check:
9^2 = 4(9)+45
81 = 36+45
81=81 (Correct!!!)